OpenStudy (anonymous):

What is the value of the expression –3x^3 when x = –3?

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

-3x^3 = -3 * -3^3 = -3 * 27 = -81 Jpenrod is correct :)